In the competitive market for organic corn, market demand is QD = 340 – 2P and market supply is



QS = 100 + 4P, where P is the price per bushel, and Q is market output in thousands of bushels.



Each individual farmer faces a marginal cost function of MC = 10 + 3q, where q is the single



farmer’s output level in thousands.




a) What is the equation for the demand (which is also MR) faced by the individual farmer?




1
Expert's answer
2022-04-05T10:32:53-0400

QD = 340 – 2P

QS = 100 + 4P

At Equilibrium, Qs=Qd

340-2P= 100+4P

6P= 240

P= 40

Q= 100+4(40)= 260